If y=|tanx|, dy/dx=____ ? Kindly answer fast....its urgent..

Grade:12

1 Answers

Arun
6 years ago
Here we have to give answer in breaks as[0,pi/2), y= tanx hence dy/dx= sec^2x(pi/2,pi], y= -tanx hence dy/dx= -sec^2x[pi,3pi/2), y= tanx hence dy/dx= sec^2x(3pi/2,2pi], y= -tanx hence dy/dx= -sec^2x